OpenStudy (thats_kyy):

A system of equations is shown below: Equation A: 2c = d − 8 Equation B: c = 3d + 8 Which of the following steps should be performed to eliminate variable d first? Multiply equation B by 2. Multiply equation A by 2. Multiply equation B by 3. Multiply equation A by −3.

OpenStudy (thats_kyy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the coefficient of the d term in equation 2?

jimthompson5910 (jim_thompson5910):

in other words, what is the number in front of the d term in equation 2?

OpenStudy (thats_kyy):

2

jimthompson5910 (jim_thompson5910):

look again

OpenStudy (thats_kyy):

2c

jimthompson5910 (jim_thompson5910):

c = 3d + 8 what is the number in front of the d term

OpenStudy (thats_kyy):

3

jimthompson5910 (jim_thompson5910):

2c = d - 8 is the same as 2c = 1d - 8 what is the number in front of the d term

OpenStudy (thats_kyy):

1

jimthompson5910 (jim_thompson5910):

we have 1d in equation 1 we have 3d in equation 2 we want the d terms to add to 0, do they?

jimthompson5910 (jim_thompson5910):

in other words, is 1d + 3d equal to 0d = 0 ?

OpenStudy (thomas5267):

Wouldn't it be quicker to directly substitute c into equation A?

OpenStudy (usukidoll):

@thomas5267 I got into a brawl with that question.

OpenStudy (triciaal):

@thomas5267 yes but the directions specifically and simply ask what should you do to eliminate d first ? not even asking to solve

OpenStudy (thomas5267):

Yes that's just a personal rant lol.

OpenStudy (triciaal):

multiply equation A by -3

OpenStudy (usukidoll):

thank you @triciaal for ending this two month old question. multiply equation A by -3 to get rid of the d is the right way to do this.